# Consistent-1-to-3: Consistent Image to 3D View Synthesis via Geometry-aware Diffusion Models

### v1

@@ 3.1 Geometry-guided Novel View Synthesis
We decompose the NVS task into two sub-tasks: (i) photometric warping, which is guided by epipolar constraints to ensure accurate alignment of observed regions, and (ii) hallucination of unseen regions using information obtained from the observed regions. To accomplish these sub-tasks, we utilize a scene representation transformer for photometric warping and a view-conditioned diffusion model for the generation of unseen regions.
@@
To render a novel image of the scene represented by $\boldsymbol{z}$, the decoder of SRT queries the pixel color via a cross attention between the ray corresponding to the pixel $\boldsymbol{r}\equiv(\boldsymbol{o},\boldsymbol{d})$ with $\boldsymbol{o}$ and $\boldsymbol{d}$ being the origin and normalized direction of the ray and the set-latent scene representation $\boldsymbol{z}$:
@@
Several changes are necessary to integrate SRT to our pipeline. Specifically, we replace all standard attention layers with epipolar attention, which allows the model to incorporate important geometry biases (details explained below). Furthermore, we train the SRT in the image latent space, so that it not only reduces computation costs but also ensures seamless compatibility with the latent diffusion model.
@@
View-conditioned Diffusion. While SRT is able to learn geometry correspondence via the attention mechanism, the pixel-level reconstruction loss leads to averaging out fine-grained details in the images, thus often resulting in blurred predictions. To capture the probabilistic nature of novel view synthesis (for the unseen parts), we employ a view-conditioned diffusion model to estimate the conditional distribution of the target view given the source view and the relative camera pose: $p\left(\boldsymbol{x}^{\mathrm{target}}\mid\boldsymbol{\pi}^{\mathrm{target}},\boldsymbol{x}^{\mathrm{source}},\boldsymbol{\pi}^{\mathrm{source}}\right)$. Note that although we follow LDM [33] to perform diffusion process in the image latent space, we still denote image in latent space with $\boldsymbol{x}$ for the simplicity.
@@
Figure 3: Details of UNet Block. After image features are extracted by the CNN, they are aggregated via the multiview attention mechanism in the transformer. The output images are conditioned on 1) the multiview feature; 2) the quried ray; and 3) the affinity matrix reweighted by the epipolar attention.
Epipolar Attention. While previous works on view-conditioned diffusion model [18] employ full attention in their model, we recognize the potential of exploiting valuable information between images through relative camera poses using epipolar constraints. To this end, we replace all cross-attention layers in our SRT and diffusion model between source and target views with our novel epipolar attention mechanism. In our implementation of epipolar attention, we reweigh the affinity matrix $A_{i,j}$ between the $i^{th}$ and $j^{th}$ view by taking the epipolar distance into account. For each pixel in view $i$, we compute the epipolar line in view $j$ and determine the epipolar distance for all pixels in view $j$. Subsequently, we convert the inverse epipolar distance into a weight map $K_{i,j}$. The affinity matrix is then reweighted using $A_{i,j}^{\prime}=A_{i,j}+K_{i,j}$.
To generate smooth Novel View Synthesis (NVS) images, it is essential to ensure not only the consistency between the generated views and the input view but also the consistency among the generated views themselves. While the novel pipeline mentioned earlier primarily emphasizes the consistency between generated views and the input view, we introduce additional beneficial techniques to enhance the consistency among the generated views. These techniques are designed to further improve the quality and coherency of the synthesized views, providing more visually pleasing and accurate images.
Multi-view Attention. In contrast to previous methods, such as Zero123 [18], which generate multiple images of the same object instance sequentially, we have observed that this sequential approach can introduce inconsistencies between generated views due to randomness in the multiple backward passes of the diffusion model. To address this issue, we propose an alternative approach where we model the joint distribution of multi-view images. This means predicting several novel views simultaneously given the source image $\boldsymbol{x}^{\mathrm{source}}$. To achieve this, as demonstrated in  Fig. 3, we modify the UNet architecture by feeding a batch of noisy multi-view images during the forward pass and extending the self-attention block to work with multi-view images such that information of different viewpoints are aggregated to produce consistent rendered images.
@@
While consistent NVS is great for 3D assets preview and many other applications, many scenarios within AR/VR still requires actual generation of 3D object for occlusion/lighting reasoning that attribute to realistic and immersive experience.
Following Zero123 [18], we adopt an optimization-based method to generate 3D assets using Score Distillation Sampling (SDS) [29]. The fundamental concept behind SDS optimization involves updating the 3D representation, often in the form of a NeRF model, by maximizing the likelihood of rendered views using a trained 2D diffusion model. When applying our model to the SDS optimization process, we sample a set of random camera poses instead of single one to accommodate the multi-view attention mentioned earlier.
@@
Figure 5: Qualitative results on NVS Consistency. While SRT (the first row) can generate views that are geometrically consistent with the input image, the rendered images are blurry. Previous diffusion-based approaches (the second and third row) are able to generate sharp and realistic-looking images, but the diffusion process leads to inconsistency. Our geometry-guided approach get the best of both worlds, generating not only plausible but also consistent images.
@@ 4.1 Experimental Setting
Dataset. We fine-tune our diffusion model on the recently released Objaverse [6] dataset , which is a large-scale CAD dataset containing 800K high-quality objects. We directly employ the processed rendering data from Zero123, which provides 12 random views for each object. For evaluation, we use the test split of Objaverse provided by Zero123. In addition, to test our model’s performance on the out-of-distribution data, we also evaluate on the Google Scanned Objects [9] dataset, which contains high-quality scanned household items. Images in all dataset are resized to $256\times 256$. The background is set to white.
Metrics. Following previous works [18], we use Peak signal-to-noise ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) to measure the similarity between rendered images and ground truth images. Moreover, we use the flow warping error $E_{\mathrm{warp}}$ [15] to quantify the consistency across different views. Specifically, we generate 32-frames video along a pre-defined circle camera trajectory for each objects, and then use the RAFT model [41] to compute the optical flow between two consecutive generated frames. The the error $E_{\mathrm{warp}}$ is defined as
@@
Baselines. We mainly evaluate our approach against methods that can generalize to open-set categories and accept single-view RGB images as inputs. In particular, we measure our performance against Zero123 [18], the first work that fine-tune view-conditioned diffusion model on Objaverse, showcasing impressive zero-shot ability. Following Zero123, we also report results of Image Variation [1], a Stable Diffusion model conditioned on images, which is also the pre-trained model we fine-tuned from. Besides, we adapt Pose-Guided Diffusion [42], a recent diffusion model for generating consistent videos using a epipolar attention mechanism similar to ours. Since the code wasn’t public, we re-implement and re-train it on Objaverse. We also add camera embedding same as ours to the model for more precise view control.
Implementation Details. Both SRT encoder and decoder are adapted from ViT-Base [8]. In encoder, the attention layers has 12 heads and 4 layers, the decoder’s attention layers consist of 6 heads and 3 layers. The ray for each patch image is parameterized in Plucker coordinates [39] and the dimension of ray embedding is 120. The MLPs in SRT have 1 hidden layer and GELU activations [12]. The VAE modules in diffusion models are freezed during training, the downsample scale of VAE is 8, which results in $32\times 32$ latent images. The basic UNet module is borrowed from Stable Diffusion. We use AdamW [20] with a learning rate of $5\times 10^{-5}$ for training. The model is fine-tuned with a batch size of 1024 on 32$\times$×A100-80GB for 7 days.
@@
ImageVariation [1] 6.42 0.506 0.578 2.24
Zero-1-to-3 [18] 18.13 0.828 0.157 1.42
$*{}^{*}$ [42] 18.92 0.841 0.141 0.70
Ours 20.72 0.877 0.112 0.37
Table 2: Quantitative results on Objaverse. We evaluate our method on the test split of Objaverse and demonstrate singificant improvements across all metrics. $*{}^{*}$ is our re-implemented version. $E_{\mathrm{warp}}\times 10^{-1}$.
@@
ImageVariation [1] 5.93 0.538 0.529 1.99
Zero-1-to-3 [18] 16.72 0.721 0.212 1.17
$*{}^{*}$ [42] 17.52 0.791 0.174 0.52
Ours 19.46 0.858 0.146 0.33
Table 3: Quantitative results on Google Scanned Objects. Evaluation of novel-view synthesis on the out-of-distribution GSO dataset. Ours is best across datasets and metrics. $E_{\mathrm{warp}}\times 10^{-1}$.
@@ 4.2 Novel View Synthesis
We present quantitative results for NVS on both in-distribution and out-of-distribution datasets in Tab. 2 and  3. Our proposed mechanisms significantly improve the performance over all metrics. Fig. 4 shows a comparison of NVS results on Objaverse with our method and all baselines. The SRT [34] results in figure is obtained by training the SRT part in our model with only the reconstruction loss. On one side, Point-E and SRT fail to produce realistic images but remain consistent with inputs; on the other side, diffusion-based baselines (Zero123 and PoseDiff) can generate realistic images but don’t align with the ground truth. Our approach is the only one achieves both fidelity and consistency.
Fig. 5 provides the qualitative results for the NVS consistency. We render a sequence of images using the same camera trajectory for each approach. While previous diffusion models (Zero123 and PoseDiff) are able to generate realistic novel views, they often suffer from the inconsistency issue across the independently sampled views. By incorporating epipolar constraint and modeling the joint distribution of multi views, our method significantly alleviates this issue.
By modeling the data distribution, diffusion approaches, by nature, can generate diverse data samples. But when dealing with NVS tasks, it’s crucial that observed regions in target views reflect the geometric transformation rather than being randomly generated. Our design is guided by this motivation, and Fig. 6 shows the effectiveness of our design. The first view of each object is close to the input view, and all generated images align with the input. But for the unseen regions shown in the second view, the variation of generated images is much larger then first view.
Figure 6: Diversity of NVS results. Nearby views (the 1st view) maintains consistency with the input image, while the distant views (the 2nd view) generate a variety of expressive yet realistic images. Note the diverse patterns on the back of the trailer.
Figure 7: 3D reconstruction. By combining the view-conditioned diffusion model and NeRF-style reconstruction techniques, our method is able to preform high-fidelity 3D reconstruction for any single images.
@@
Fig. 7 qualitatively shows the 3D reconstruction results of our method. For each instance, we optimize a NeRF via the SDS loss, the colored images in the figure are all NeRF renderings. After reconstructing NeRFs, we further utilize DiffRast [16] to extract meshes from NeRFs. In this way, we leverage the multi-view prior from our diffusion models and combines them with the advanced NeRF-style reconstruction.
@@
w.o. epi attention 19.34 0.847 0.148 0.43
w.o. in-out attention 18.79 0.823 0.156 0.46
w.o. mv-out attention 20.84 0.879 0.120 0.70
w.o. hierarchy generation 21.03 0.883 0.107 0.88
full model 20.72 0.877 0.112 0.37
@@ 4.4 Ablation Study
We ablate each design choices in our method and present quantitative results in Tab. 4. The epipolar line guidance (1st row) and the attention between inputs and outputs (2nd row) enforce the model to learn better geometric transformation of observed regions in source views, and thus they are essential to the similarity metrics (PSNR, SSIM and LPIPS). The attention between multiple outputs (3rd row) and the hierarchy generation (4th row) contributes a lot to the consistency across different views. However, it should be noted that these two mechanisms will slightly hurt the reconstruction performance, and our full model is actually a trade-off between fidelity and consistency.
@@ 5 Conclusion
In this paper, we proposed Consistent 1-to-3, a geometry-aware deep model that produce high quality and 3D consistent novel view synthesis from a single image or few-short images. Consistent 1-to-3 is a two stage model that first uses a transformer to generate a blurry while geometry correct image, and then uses a diffusion model to finish the details. We demonstrate by following such a generation progression, the generated results are significantly improved. We hope our work could motivate future studies on incorporating better geometry constraints and representations inside the model, which could significantly improve its practical usage.

### v2

@@ 3.1 Geometry-guided Novel View Synthesis
We decompose the NVS task into two sub-tasks: (i) photometric warping, which is guided by epipolar constraints to ensure accurate alignment of observed regions, and (ii) hallucination of unseen regions based on observed regions. To accomplish these sub-tasks, we utilize a scene representation transformer for photometric warping and a view-conditioned diffusion model for the generation of unseen regions.
@@
To render a novel image of the scene represented by $\boldsymbol{z}$, the decoder of SRT queries the pixel color via cross attention between the ray corresponding to the pixel $\boldsymbol{r}\equiv(\boldsymbol{o},\boldsymbol{d})$ with $\boldsymbol{o}$ and $\boldsymbol{d}$ being the origin and normalized direction of the ray and the set-latent scene representation $\boldsymbol{z}$:
@@
Several changes are necessary to integrate SRT into our pipeline. Specifically, we replace all standard attention layers with epipolar attention, which allows the model to incorporate important geometry biases (details explained below). Furthermore, we train the SRT in the image latent space, so that it not only reduces computation costs but also ensures seamless compatibility with the latent diffusion model.
@@
View-conditioned Diffusion. While SRT is able to learn geometry correspondence via the attention mechanism, the pixel-level reconstruction loss leads to averaging out fine-grained details in the images, thus often resulting in blurred predictions. To capture the probabilistic nature of novel view synthesis (for the unseen parts), we employ a view-conditioned diffusion model to estimate the conditional distribution of the target view given the source view and the relative camera pose: $p\left(\boldsymbol{x}^{\mathrm{target}}\mid\boldsymbol{\pi}^{\mathrm{target}},\boldsymbol{x}^{\mathrm{source}},\boldsymbol{\pi}^{\mathrm{source}}\right)$. Note that although we follow LDM [37] to perform the diffusion process in the image latent space, we still denote the image in latent space with $\boldsymbol{x}$ for simplicity.
@@
Figure 3: Details of UNet Block. After image features are extracted by CNN, they are aggregated via the multi-view attention mechanism in the transformer. The output images are conditioned on 1) the multi-view feature; 2) the queried ray; and 3) the affinity matrix reweighted by the epipolar attention.
Epipolar Attention. While previous works on the view-conditioned diffusion model [22] employ full attention in their model, we propose to exploit the valuable information between images through relative camera poses using epipolar constraints. To this end, we replace all cross-attention layers in our SRT and diffusion model between source and target views with our novel epipolar attention mechanism. In our implementation of epipolar attention, we reweigh the affinity matrix $A_{i,j}$ between the $i^{th}$ and $j^{th}$ view by taking the epipolar distance into account. For each pixel in view $i$, we compute the epipolar line in view $j$ and determine the epipolar distance for all pixels in view $j$. Subsequently, we convert the inverse epipolar distance into a weight map $K_{i,j}$. The affinity matrix is then reweighted using $A_{i,j}^{\prime}=A_{i,j}+K_{i,j}$.
3.2 Improving Multi-view Consistency
To generate smooth Novel View Synthesis (NVS) images, it is essential to ensure not only the consistency between the generated views and input view but also the consistency among the generated views themselves. While our novel pipeline mentioned earlier primarily emphasizes the consistency between generated and input views, we introduce additional beneficial techniques to enhance the consistency among the generated views. These techniques are designed to further improve the quality and coherency of the synthesized views, providing more visually pleasing and accurate images.
Multi-view Attention. In contrast to previous methods, such as Zero123 [22], which generate multiple images of a givens object instance sequentially, we have observed that this sequential approach can introduce inconsistencies between generated views due to randomness in the multiple backward passes of the diffusion model. To address this issue, we propose an alternative approach to learn the joint distribution of multi-view images. This means predicting several novel views simultaneously given the source image $\boldsymbol{x}^{\mathrm{source}}$. To achieve this, as demonstrated in  Fig. 3, we modify the UNet architecture by feeding a batch of noisy multi-view images during the forward pass and extending the self-attention block to work with multi-view images such that information of different viewpoints is aggregated to produce consistent rendered images.
@@ 3.2 Improving Multi-view Consistency
While consistent NVS is great for 3D assets preview and many other applications, many scenarios within AR/VR still require accurate 3D objects for occlusion/lighting reasoning that is attributed to realistic and immersive experience.
Following Zero123 [22], we adopt an optimization-based method to generate 3D assets using Score Distillation Sampling (SDS) [33]. The fundamental concept behind SDS optimization involves updating the 3D representation, often in the form of a NeRF model, by maximizing the likelihood of rendered views using a trained 2D diffusion model. When applying our model to the SDS optimization process, we sample a set of random camera poses instead of a single one to accommodate the multi-view attention mentioned earlier.
@@
Figure 5: Qualitative results on NVS Consistency. While SRT (the first row) can generate views that are geometrically consistent with the input image, the rendered images are blurry. Previous diffusion-based approaches (the second and third row) are able to generate sharp and realistic-looking images, but the diffusion process leads to inconsistency. Our geometry-guided approach gets the best of both worlds, generating not only plausible but also consistent images.
@@ 4.1 Experimental Setting
Dataset. We fine-tune our diffusion model on the recently released Objaverse [7] dataset, which is a large-scale CAD dataset containing 800K high-quality objects. We directly employ the processed rendering data from Zero123, which provides 12 random views for each object. For evaluation, we use the test split of Objaverse provided by Zero123. In addition, to test our model’s performance on the out-of-distribution data, we also evaluate on the Google Scanned Objects [10] dataset, which contains high-quality scanned household items. Images in all datasets are resized to $256\times 256$. The background is set to white.
Metrics. Following previous works [22], we use Peak signal-to-noise ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) to measure the similarity between rendered images and ground truth images. Moreover, we use the flow warping error $E_{\mathrm{warp}}$ [17] to quantify the consistency across different views. Specifically, we generate a 32-frame video along a pre-defined circle of camera trajectory for each object and then use the RAFT model [47] to compute the optical flow between two consecutive generated frames. The error $E_{\mathrm{warp}}$ is defined as
@@
Baselines. We mainly evaluate our approach against methods that can generalize to open-set categories and accept single-view RGB images as inputs. In particular, we measure our performance against Zero123 [22], the first work that fine-tune the view-conditioned diffusion model on Objaverse, showcasing impressive zero-shot ability. Following Zero123, we also report results of Image Variation [1], a Stable Diffusion model conditioned on images, which is also the pre-trained model we fine-tuned from. Besides, we adapt Pose-Guided Diffusion [48], a recent diffusion model for generating consistent videos using an epipolar attention mechanism similar to ours. Since the code wasn’t public, we re-implement and re-train it on Objaverse. We also add a camera embedding the same as ours to the model for more precise view control.
Implementation Details. Both SRT encoder and decoder are adapted from ViT-Base [9]. In encoder, the attention layers has 12 heads and 4 layers, the decoder’s attention layers consist of 6 heads and 3 layers. The ray for each image patch is parameterized in Plucker coordinates [45] and the dimension of ray embedding is 120. The MLPs in SRT have 1 hidden layer and GELU activations [14]. The VAE modules in diffusion models are freezed during training. The downsample scale of VAE is 8, which results in $32\times 32$ latent images. The basic UNet module is borrowed from Stable Diffusion. We use AdamW [24] with a learning rate of $5\times 10^{-5}$ for training. The model is fine-tuned with a batch size of 1024 on 32$\times$×A100-80GB for 7 days.
@@
ImageVariation [1] 6.42 0.506 0.578 22.4
Zero-1-to-3 [22] 18.13 0.828 0.157 14.2
$*{}^{*}$ [48] 18.92 0.841 0.141 7.0
Ours 20.72 0.877 0.112 3.7
Table 2: Quantitative results on Objaverse. We evaluate our method on the test split of Objaverse and demonstrate significant improvements across all metrics. $*{}^{*}$ is our re-implemented version.
@@
ImageVariation [1] 5.93 0.538 0.529 19.9
Zero-1-to-3 [22] 16.72 0.721 0.212 11.7
$*{}^{*}$ [48] 17.52 0.791 0.174 5.2
Ours 19.46 0.858 0.146 3.3
Table 3: Quantitative results on Google Scanned Objects. Evaluation of novel-view synthesis on the out-of-distribution GSO dataset. Ours is best across datasets and metrics.
@@ 4.2 Novel View Synthesis
We present quantitative results for NVS on both in-distribution and out-of-distribution datasets in Tab. 2 and  3. Our proposed mechanisms significantly improve the performance of all metrics. Fig. 4 shows a comparison of NVS results on Objaverse with our method and all baselines. The SRT [39] results are obtained by training the SRT part of our model with only the reconstruction loss. On one side, Point-E and SRT fail to produce realistic images but remain consistent with the input; on the other side, diffusion-based baselines (Zero123 and PoseDiff) can generate realistic images but do not align well with the ground truth. Our approach is the only one that achieves both fidelity and consistency.
Fig. 5 provides the qualitative results for the NVS consistency. We render a sequence of images using the same camera trajectory for each approach. While previous diffusion models (Zero123 and PoseDiff) are able to generate realistic novel views, they often suffer from inconsistent issues across the independently sampled views. By incorporating epipolar constraint and modeling the joint distribution of multi-views, our method significantly alleviates this issue.
By learning the data distribution, diffusion approaches, by nature, can generate more diverse samples. But when dealing with NVS tasks, observed regions in target views must reflect the geometric transformation rather than being randomly generated. Our design is guided by this motivation, and Fig. 6 shows the effectiveness of our design. The first view of each object is close to the input view, and all generated images align with the input. But for the unseen regions shown in the second view, the variation of generated images is much larger the first view.
Figure 6: Diversity of NVS results. Nearby views (the 1st view) maintain consistency with the input image, while distant views (the 2nd view) generate a variety of expressive yet realistic images. Note the diverse patterns on the back of the trailer.
Figure 7: 3D reconstruction. By combining the view-conditioned diffusion model and NeRF-style reconstruction techniques, our method is able to perform high-fidelity 3D reconstruction for any single image.
@@
Fig. 7 qualitatively shows the 3D reconstruction results of our method. For each instance, we optimize a NeRF via the SDS loss, the colored images in the figure are all NeRF renderings. After reconstructing NeRFs, we further utilize DiffRast [18] to extract meshes from NeRFs. In this way, we leverage the multi-view prior from our diffusion models and combine them with the advanced NeRF-style reconstruction.
@@
w.o. epi attention 19.34 0.847 0.148 4.3
w.o. in-out attention 18.79 0.823 0.156 4.6
w.o. mv-out attention 20.84 0.879 0.120 7.0
w.o. hierarchy generation 21.03 0.883 0.107 8.8
full model 20.72 0.877 0.112 3.7
@@ 4.4 Ablation Study
We ablate each design choice in our method and present quantitative results in Tab. 4. The epipolar line guidance (1st row) and the attention between inputs and outputs (2nd row) enforce the model to learn the better geometric transformation of observed regions in source views, and thus they are essential to the similarity metrics (PSNR, SSIM, and LPIPS). The attention between multiple outputs (3rd row) and the hierarchy generation (4th row) contributes a lot to the consistency across different views. However, it should be noted that these two mechanisms will slightly hurt the reconstruction performance, and our full model is a trade-off between fidelity and consistency.
@@ 5 Conclusion
In this paper, we proposed Consistent 1-to-3, a geometry-aware deep model that produces high quality and 3D consistent novel view synthesis from a single image or few-shot images. Consistent 1-to-3 is a two-stage model that first uses a transformer to generate a blurry while geometry correct image, and then uses a diffusion model to finish the details. We demonstrate by following such a generation progression, the generated results are significantly improved. We hope our work could motivate future studies on incorporating better geometry constraints and representations inside the model, which could significantly improve its practical usage.
Supplementary Material
6 Overview
In this supplementary material, we first provide the technical details of the proposed method to complement the paper. Second, we describe the experiment details, including the metric computation and reproducing baselines. Finally, we present additional qualitative results of baselines and our approach. Please see our video for an overview of our work and more visualizations.
7 Technical Details
Epipolar Attention. The epipolar line computation process is visualized in Fig. 8 (a). Given a point $\boldsymbol{p}$ on the image plane of the target view $i$ and the relative camera pose $\boldsymbol{\pi}^{i\to j}$, the goal is the compute the corresponding epipolar line $\boldsymbol{e}$ on the image plane of the source view $j$. We first compute the 3D ray of point $\boldsymbol{p}$, then we project any two 3D points on this ray onto the source view image plane. The line connected by these two projected points is the epipolar line $\boldsymbol{e}$.
Figure 8: By combining the view-conditioned diffusion model and NeRF-style reconstruction techniques, our method is able to preform high-fidelity 3D reconstruction for any single images.
As mentioned in the main text, after getting the epipolr line, we compute pixel distances between patch centers from different views and invert it to weight map $K_{i,j}$. Finally, we directly add $K_{i,j}$ to the original attention weights $A_{i,j}$: $A^{\prime}_{i,j}=A_{i,j}+K_{i,j}$. In Fig. 8 (b), for the blue point in the left figure, we show its corresponding inverted epipolar distances on the right side.
UNet Block. The UNet block consists of three modules: (i) a convolution layer; (ii) a self-attention layer among multiple output views; (iii) Cross-attention layer between input views and output views. The noisy latent images $\boldsymbol{y}$ (concatenated with $\tilde{\boldsymbol{x}}^{\mathrm{target}}$ predicted by SRT) are first passed to a convolution layer to obtain patch-level image features. Secondly, instead of performing self-attention within each individual view, we merge tokens from multiple output views and apply self-attention across all these tokens. Finally, we concatenate patch tokens with the corresponding ray embedding, and apply an epipolar attention between individual output views and input tokens (obtained by SRT).
Hyper-parameters. Hyper-parameters used in our method are reported in Tab. 5.
Parameter Value
Input view num 1 - 4
Output view num 1 - 4
Image size 256
Latent image size (after vae) 32
SRT feature dim 768
SRT head num 12
SRT activation gelu
SRT-Enc layer num 8
SRT-Dec layer num 4
UNet down/up block num 4
UNet block out channels (320, 640, 1280, 1280)
UNet in channels 8
UNet out channels 4
UNet activation silu
Diffusion steps (training) 1000
Diffusion steps (inference) 50
Sampling method PNDM
Optimizier AdamW
Learning rate $5e^{-5}$
AdamW Betas (0.9, 0.999)
Iterations 100,000
Batch Size 1024
Diffusion loss weight 1.0
Reconstruction loss weight 0.5
Table 5: We report main hyper-parameters used in our method here.
8 Experiment Details
Dataset. We directly use the rendering data and the training/test split of the Objaverse dataset [7] provided by Zero123 [22]. The whole dataset contains 800k objects, with 7k for testing. For each object, Zero123 provides 12 random views sampled from a uniform sphere. In addition, we employ the Google Scanned Object (GSO) dataset to evaluate the zero-shot ability of our method. For GSO, we directly use the rendering data provided by IBRNet [51]. During evaluation, we take the first frame of each object as the source frame and employ the subsequent frame as the target frame.
Metrics. We employ the VGG model, pre-trained on the ImageNet dataset, to calculate the LPIPS score. While computing the flow warping error $E_{\mathrm{warp}}$, the pre-trained RAFT [47] model is used to to compute the optical flow between consequent frames. For each object, a 32-frame video is generated where phi spans 360 degree, and the elevation is set to 10 degree with a radius of 1.5.
Reproducing Baselines.
  • Zero123. We use the official implementation on GitHub [^1] https://github.com/cvlab-columbia/zero123 and the checkpoint pre-trained on Objaverse. We keep all the hyper-parameters and use their demo example to perform novel view synthesis. During inference, the guidance scale is set to 3.0 and the image is resized to $256\times 256$.
  • Point-E. We use the official implementation and pre-trained models on GitHub [^2] https://github.com/openai/point-e. We follow their example demo to perform 3D reconstruction.
  • PoseDiff. While the code is publicly available at the moment, we faithfully re-implement their model based on the description in the paper. All hyper-parameters are consistent with our approach.